OpenStudy (anonymous):

What is the equation of the line that contains the points with (x,y) coordinates (-3, 7) and (5, -1)? a. y=3x-2 b. y=x+10 c. y=-x +4 d. y= - (1/3)x +8 e. y= -(3/2)x + (11/4)

OpenStudy (anonymous):

slope=(7+1)/(-3-5) =8/-8=-1 (y+1)/(x-5)=-1 y+1=5-x x+y=4 y=-x+4

OpenStudy (anonymous):

here we take any point(x,y) on line.